An electronic device designed to operate at any temperature in the range from -5°C to 72 °C contains pure carbon resistors. By what factor does their resistance increase over this range? Let the initial temperature be 23 °C.


Round your answer to 3 decimal places


1
Expert's answer
2021-05-14T18:23:16-0400
"\frac{R}{R_0}=1+\alpha \Delta T\\ \frac{R}{R_0}=1+(10^{-5})(72-23)=1.0005"